# UTTERTUNE: LORA-BASED TARGET-LANGUAGE PRONUNCIATION EDIT AND CONTROL IN MULTILINGUAL TEXT-TO-SPEECH

### v1

@@ 1 Introduction
Text-to-speech (TTS) models based on large language model (LLM) architecture (LLM-TTS in this paper) have demonstrated exceptional naturalness, especially in zero-shot multi-speaker and multilingual synthesis, leading the way in speech synthesis technology [1, 2, 3, 4]; however, reproducing accurate pronunciation remains challenging. Some multilingual LLM-TTS, such as CosyVoice 2 [4], are designed to take raw text (characters) as input and tokenize it via byte-pair encoding (BPE) [5], without explicit phonemic or prosody markers. This design contrasts with conventional neural sequence-to-sequence TTS, which typically converts input text into phonemes (grapheme-to-phoneme; G2P) and prosody information, if needed, using a text frontend before feeding it into the model [6, 7, 8, 1]. Omitting the text frontend simplifies model architecture and facilitates multilingual training. On the other hand, such models require a large amount of speech-text pairs that cover the diversity of the target language because they predict segmental pronunciation and prosody data-driven. This dependency engenders known difficulties with some languages.
@@ 4.1 Training and Validation Data
The training process thus pairs these transcripts with the original speech recordings. As a result, the dataset comprised approximately 15,097 Japanese speech-transcription pairs. We held out 754 utterances from it for validation.
@@ 4.2 Evaluation Data
Test Set 2 consisted of 50 sentences designed to assess the controllability of segmental pronunciation and pitch accent. The sentences were obtained by instructing ChatGPT o3 to generate easy-to-read Japanese sentences that include one difficult-to-read word each because no appropriate public benchmarks existed. After the generation, the author also reviewed all the generated sentences and made necessary modifications. The difficult words lasted from 2 to 6 morae, and the accent nuclei were located from the first to the last morae. For the proposed method, the pronunciation of the target difficult-to-read word was spelled out in the <PHON_START><PHON_END> format. We performed zero-shot synthesis on the same speakers used in Test Set 1. As a result, we obtained 2,100 stimuli per system.
@@ 4.5 Evaluation Metrics
CosyVoice 2-0.5B (kana input) 3.28 - 0.692 0.0595 0.472
Proposed 3.30 - 0.697 0.0626 0.975
@@
We performed a subjective accent evaluation on Test Set 2. In this paper, due to time constraints, the author carefully listened to each synthesized speech and judged whether the pitch accent pattern of the target word followed the instruction. Since this metric required highly specialized manual work, we limited the speakers to four out of 42. The four speakers were chosen as those closest to the centroids of four clusters formed by k-means clustering, using uniform manifold approximation and projection [35] to project ERes2Net embeddings (192 dimensions) extracted from the prompt audio of the 42 speakers onto two dimensions.
@@ 5.2 Effect of Phoneme Tags and Kana Transcription
UtterTune, by contrast, attained 0.975 accent correctness while keeping a competitive CER (0.0626). The difference of CER to the kana-baseline was significant ($p<0.005$), but the effect size was negligible (Cliff’s delta=0.115=0.115 [36]). The LoRA adaptation learnt not only to obey the given phoneme string but also to place the pitch fall at the right mora.
To confirm that the effect was not extended to the parts not enclosed by <PHON_START> and <PHON_END>, an additional 50 sentences containing only one katakana word with an accented nucleus were prepared based on the ChatGPT o3 generation. If affected, these words should be synthesized with a flat pitch pattern. The sentences were synthesized by the four speakers using the baseline and the proposed method. The percentage of correct accents, scored by the author, was 0.925 (baseline) and 0.905 (proposed method), with no significant difference, and a 99% CI of [-0.080, 0.045]. This means that UtterTune did not leak pronunciation controllability outside the tag tokens.
@@ 6 Conclusion
We proposed UtterTune, a lightweight LoRA-based adaptation that endows a multilingual LLM-TTS model with phoneme-level pronunciation control. By freezing the 0.5-B-parameter CosyVoice 2 backbone and fine‑tuning only rank‑16 LoRA matrices plus two special input tokens, UtterTune requires less than 0.5% additional parameters and a few GPU‑hours of training. It raised naturalness from 3.44 to 3.88 MOS, kept CER below 7% on everyday sentences, and boosted accent correctness from 0.472 to 0.975 on stress‑test material without degrading speaker similarity. UtterTune never deteriorates in the performance of other languages in principle because LoRA does not need to be applied when synthesizing other languages. These results demonstrate that minor, targeted updates can rectify language-specific deficiencies—here, Japanese G2P and pitch accent—while preserving the strong performance of a multilingual model.

### v2

@@ 1 Introduction
Text-to-speech (TTS) models based on large language model (LLM) architecture (LLM-TTS in this paper) have demonstrated exceptional naturalness, especially in zero-shot multi-speaker and multilingual synthesis, leading the way in speech synthesis technology [1, 2, 3, 4]; however, reproducing accurate pronunciation remains challenging. Some multilingual LLM-TTS, such as CosyVoice 2 [4], are designed to take raw text (characters) as input and tokenize it via byte-pair encoding (BPE) [5], without explicit phonemic or prosody markers. This design contrasts with conventional neural sequence-to-sequence TTS, which typically converts input text into phonemes (grapheme-to-phoneme; G2P) and prosody information, if needed, using a text frontend before feeding it into the model [6, 7, 8, 1]. Omitting the text frontend simplifies model architecture and facilitates multilingual training. On the other hand, such models require a large amount of speech-text pairs that cover the diversity of the target language because they predict segmental pronunciation and prosody data-driven. This dependency gives rise to known difficulties with some languages.
@@ 4.1 Training and Validation Data
The training process thus pairs these transcripts with the original speech recordings. As a result, the dataset comprised 15,097 Japanese speech-transcription pairs. We held out 754 utterances from it for validation.
@@ 4.2 Evaluation Data
Test Set 2 consisted of 50 sentences designed to assess the controllability of segmental pronunciation and pitch accent. The sentences were obtained by instructing ChatGPT o3 to generate easy-to-read Japanese sentences that include one difficult-to-read word each because no appropriate public benchmarks existed. After the generation, the author also reviewed all the generated sentences and made necessary modifications. The difficult words lasted from 2 to 6 morae, and the accent nuclei were located from the first to the last morae. Accent annotations followed the NHK Japanese Accent Dictionary [11] for the words where the entry existed. For the proposed method, the pronunciation of the target difficult-to-read word was spelled out in the <PHON_START><PHON_END> format. We performed zero-shot synthesis on the same speakers used in Test Set 1. As a result, we obtained 2,100 stimuli per system.
@@ 4.5 Evaluation Metrics
CosyVoice 2-0.5B (kana input) 3.28 - 0.692 0.0595 0.498
Proposed 3.30 - 0.697 0.0626 0.899
@@
We conducted another subjective listening test where participants listened to and identified whether the pitch accent of the target word in the synthesized speech was spoken as instructed or not. 15 crowdworkers who are native Japanese speakers and passed a check test [^1] It is known that even native Japanese speakers find it challenging to identify the Tokyo Japanese accent when listening accurately. participated in this test. We gathered MOS scores for 126 stimuli paired for each model, stratified by sentence number from Test Set 2, with a minimum of two ratings per sample. The stimuli were selected to contain all 42 speakers.
@@ 5.2 Effect of Phoneme Tags and Kana Transcription
UtterTune, by contrast, attained 0.899 accent correctness while keeping a competitive CER (0.0626). The difference of CER to the kana-baseline was significant ($p<0.005$), but the effect size was negligible (Cliff’s delta=0.115=0.115 [35]). The LoRA adaptation learnt not only to obey the given phoneme string but also to place the pitch fall at the right mora.
To confirm that the effect was not extended to the parts not enclosed by <PHON_START> and <PHON_END>, an additional 50 sentences containing only one katakana word with an accented nucleus were prepared based on the ChatGPT o3 generation. Accent annotations followed the NHK Japanese Accent Dictionary [11]. If affected, these words should be synthesized with a flat pitch pattern. The sentences were synthesized by all 42 speakers using the baseline and the proposed method. The percentage of correct accents, judged in the same procedure as measuring the accent correctness, was 0.796 (baseline) and 0.829 (proposed method), with no significant difference. This means that UtterTune did not leak pronunciation controllability outside the tag tokens.
@@ 6 Conclusion
We proposed UtterTune, a lightweight LoRA-based adaptation that enables a multilingual LLM-TTS model with phoneme-level pronunciation control. By freezing the 0.5-B-parameter CosyVoice 2 backbone and fine‑tuning only rank‑16 LoRA matrices plus two special input tokens, UtterTune requires less than 0.5% additional parameters and a few GPU‑hours of training. It raised naturalness from 3.44 to 3.88 MOS, kept CER below 7% on everyday sentences, and boosted accent correctness from 0.498 to 0.899 on stress‑test material without degrading speaker similarity. UtterTune never deteriorates in the performance of other languages in principle because LoRA does not need to be applied when synthesizing other languages. These results demonstrate that minor, targeted updates can rectify language-specific deficiencies—here, Japanese G2P and pitch accent—while preserving the strong performance of a multilingual model.